OpenStudy (anonymous):

Kathy wants to factor x^2-12x+36 by grouping; however, Tom says it is a special product and can factor a different way. Using complete sentences, explain and demonstrate how both methods will result in the same factors.

OpenStudy (ranga):

To factor by grouping, rewrite it as: x^2-12x+36 = x^2 - 6x - 6x + 26 Now you can factor. Alternative way: Compare x^2 - 6x + 36 to: (a - b)^2 = a^2 - 2ab + b^2 and you will recognize it is a special factor.

OpenStudy (anonymous):

Thank you :)

OpenStudy (ranga):

you are welcome.